# U-REPA: Aligning Diffusion U-Nets to ViTs

### v1

@@ Abstract
Representation Alignment (REPA) that aligns Diffusion Transformer (DiT) hidden-states with ViT visual encoders has proven highly effective in DiT training, demonstrating superior convergence properties, but it has not been validated on the canonical diffusion U-Net architecture that shows faster convergence compared to DiTs. However, adapting REPA to U-Net architectures presents unique challenges: (1) different block functionalities necessitate revised alignment strategies; (2) spatial-dimension inconsistencies emerge from U-Net’s spatial downsampling operations; (3) space gaps between U-Net and ViT hinder the effectiveness of tokenwise alignment. To encounter these challenges, we propose U-REPA, a representation alignment paradigm that bridges U-Net hidden states and ViT features as follows: Firstly, we propose via observation that due to skip connection, the middle stage of U-Net is the best alignment option. Secondly, we propose upsampling of U-Net features after passing them through MLPs. Thirdly, we observe difficulty when performing tokenwise similarity alignment, and further introduces a manifold loss that regularizes the relative similarity between samples. Experiments indicate that the resulting U-REPA could achieve excellent generation quality and greatly accelerates the convergence speed. With CFG guidance interval, U-REPA could reach in 200 epochs or 1M iterations on ImageNet 256 256, and needs only half the total epochs to perform better than REPA. Codes are available at https://github.com/YuchuanTian/U-REPA.
@@ 1 Introduction
Representation Alignment (REPA) [45], a methodology that aligns features from Diffusion Transformers (DiT) [30] to modern visual encoders, has been demonstrated to significantly accelerate DiT training. This approach holds particular significance given the growing prominence of DiTs, which have gained mainstream adoption in diffusion models and are extensively applied across image generation [4; 11; 23] and video generation domains [48; 22; 19]. However, emerging empirical evidence suggests that U-Net [33] architectures might present a more advantageous alternative to DiTs in certain scenarios [18; 7; 39; 38]: U-Net-based models exhibit substantially faster convergence while achieving generation quality comparable to their transformer-based counterparts. This dichotomy motivates our core research inquiry - can modern Vision Transformer (ViT [10])-based visual encoders be effectively adapted to guide diffusion U-Net training through alignment mechanisms similar to REPA, thereby potentially elevating the convergence speed ceiling of diffusion models?
@@ 2 Related Work
The development of diffusion architectures. The conventional diffusion works [17; 35; 36; 9] leverages a U-Net [33] architecture, whose basic block is a concatenation of convolution layers and self-attention. More recent architectural innovations in diffusion models have witnessed a paradigm shift from conventional U-Net frameworks toward transformer-based architectures. The emergence of Diffusion Transformers [1; 30] demonstrates their competitive performance despite abandoning the inductive biases inherent in U-Net designs. U-ViT [1] represents an intermediate architecture that preserves U-Net’s hierarchical structure but replaces convolutional blocks with transformer layers, notably omitting the traditional downsampling operations. Subsequent developments have further streamlined the architecture: DiT [30] adopts a pure transformer backbone with isotropic scaling, while SiT [26] integrates the transformer architecture into the RectifiedFlow framework. Some other works either improve the micro-designs [6; 25], or focuses on architectural efficiency [3; 42; 40].
In contrary to these DiT works, some works still sticks to U-Net architectures and offer valuable rethinking on this conventional architectural preference: in pixel-space image generation, works including SimpleDiffusion [18] and HourglassDiT [7] still sticks to U-Net; in with its variants like U-DiT [39], Playground v3 [24], and DiC [38] extending its success to latent-space diffusion through simple Conv33 designs. While these implementations empirically validate U-Net’s accelerated convergence and stable training dynamics compared to transformer-based alternatives, current research predominantly focuses on proposing architectural modifications rather than uncovering the reasons of U-Net’s superior diffusion performance.
Techniques for better DiT performance. Building upon the success of self-supervised learning [16], MDT [13] and MaskDiT [47] pioneer masked image modeling in diffusion frameworks by adaptively masking a good proportion of input patches during training. Other than the masking strategy, a bunch of diffusion works refer to higher-level semantic guidance from off-the-shelf pretrained models that significantly improves generation quality. REPA [45] establishes feature alignment between ViT-based encoder embeddings and diffusion latent spaces through contrastive learning. LightningDiT [44] innovates through an improved VAE distilled from MAE [16] and DINOv2 [28]. Ma et al. [27] introduces CLIP [31] and DINO [2] to verify inference-time scaling of diffusion models. These methods demonstrate that a higher-level semantic-rich feature-map from pretrained vision encoders is helpful to diffusion-based generation.
@@ 3.1 Preliminaries: REPA for DiT
Representation Alignment (REPA) [45] distills Diffusion Transformers with semantic features from off-the-shelf ViT-based vision encoders (e.g. DINOv2 [28], CLIP [31], MAE [16], et cetera). Given a ViT-based vision encoder and clean image , let denote its patch embeddings, where and represent the number of patches and embedding dimension, respectively. REPA establishes feature alignment between and the projected diffusion encoder outputs , where is the latent representation from the diffusion transformer at timestep , and is a trainable multilayer perceptron (MLP).
@@
where denotes a similarity metric (e.g., cosine similarity). Typically, is adopted as the output from early layers (the original work adopts layer index 8) in DiT for better alignment. This alignment term is combined with the basic flow-based diffusion objective (i.e. SiT [26]) through a tunable coefficient , and the final loss for diffusion model training is formulated as follows:
@@ 3.2 Evaluating the Potential of U-Net
In diffusion models, U-Net and isotropic architectures (e.g., DiT) exhibit distinct design philosophies. While DiT achieves state-of-the-art results through scalability and integration with advanced techniques, U-Net-based methods emphasize faster convergence [39]. To dissect U-Net’s efficacy, we isolate its two core components: skip connections and downsampling.
@@ 3.3 Aligning U-Net to ViT Encoders
Position for alignment. Regarding the concern that block functionalities differ between U-Net and DiT, the comparison of prior studies [12; 5] demonstrates divergent hierarchical specialization: U-Net architectures typically employ mid-network layers for high-level semantic synthesis while reserving shallow layers for low-level image refinement, whereas DiTs exhibit a totally different pattern - early layers primarily govern semantic-rich outline formation with deeper layers handling detailed image refinement.
@@
Some recent works on Diffusion U-Net [41; 34] reveals that higher-stage U-Net features are low-frequency subspaces that discards higher frequency compoenents, including noises. Gaps are inevitable when evaluating the cosine similarities of detail-rich, high-frequency-rich vectors and flat, low-frequency-dominated vectors. In this sense, strict token-level alignment constraints like the original REPA loss prove suboptimal under these conditions, as they assume implicit feature space homogeneity between aligned modalities.
@@
Rather than strict token-wise regularization, we resort to looser objectives that does not require rigid augular alignment. Inspired by manifold knowledge distillation [14], we hold that aligning similarities between samples from the same feature space could be a promising solution. Hence, we define Manifold Loss as
@@ 3.4 Other Improvements
We also propose and evaluate some other improvements as follows. Due to page limits, the corresponding ablations are enclosed in the Appendix.
Time-aware MLP. Some works on in the diffusion task [29; 46] reveals that a channel dimension is particularly sensitive and useful to a certain subset of time during sampling. As the time-variant feature is to be aligned to time-invariant vision encoder features, we hold that alignment could perform better when the MLP is time-aware and extract time-invariant information out of diffusion features for alignment.
Inspired by the conditioning of canonical Diffusion U-Net [17; 36; 9] and DiT [30], we add a module to predict a pair of channel-wise shift& scale vector . The module is in parallel to MLP and follows the design of DiT’s AdaLN, which is a concatenation of a SiLU and a Linear layer. The shift& scale vectors are imposed on the output MLP as follows:
Weight schedule of loss. As is prompted in REPA [45], designing weight schedule is a future direction. We try various weight schedules (i.e. make in Eq. 2 a function with respect to time) but found that these schedules bring very limited improvements on the proposed U-REPA. Hence, we stick to the original constant weight strategy of REPA.
@@ 4.1 Experiment Setup
Experiment settings. Our implementation completely adheres to the training protocol established in REPA [45]. Following the architectural configuration of latent diffusion models [32], we employ the identical VAE variant (sd-vae-ft-ema) and adopt the AdamW optimizer. To ensure fair comparison, we maintain identical hyperparameter settings across all experiments: a global batch size of 256, fixed learning rate of , and disabled weight decay (set to 0). () is set as (0.9, 0.999). All experiments are conducted on the ImageNet 2012 benchmark [8] under a controlled environment with a fixed random seed (global seed=0).
For main experiments (Tab. 4), we apply guidance interval [21] and SDE sampling according to the convention of REPA [45] for fair comparison. We select smaller cfg of 1.65, because we found it is better for our architecture, different from SiT. For all ablation experiments, we train models for 100K iterations, which is sufficient to show the trend of model performance; sampling is conducted with the default setting of the official REPA codebase, i.e. in ODE and guidance interval .
@@ 4.2 The Advantage of U-REPA
Comparing SiT with SiT at different scales. We evaluate our U-REPA alignment method on ImageNet 256 under a generation setting with (REPA framework without classifier-free guidance). As shown in Table 3, our approach consistently improves generation quality while significantly reducing computational costs across model scales. For the base-size SiT-B/2 variant, integrating U-REPA achieves a 39.3% improvement in FID (from 24.4 to 15.3) with comparable FLOPs (24.1G vs. 23.0G) and identical training iterations (400K), demonstrating that feature alignment enhances parameter efficiency without additional training overhead. The acceleration effect becomes more pronounced in larger models: for SiT-L/2, U-REPA reduces required iterations by 42.9% (700K→400K) while simultaneously lowering FLOPs (79.3G vs. 80.8G) and achieving a 30.9% FID improvement (8.4→5.8). Most notably, the XL-scale variant with U-REPA (cf. Fig. 4 for FIDs vs. Training iters) attains state-of-the-art FID (5.4) using 90% fewer iterations (400K vs. 4M) and fewer FLOPs (108.8G vs. 118.6G) compared to the baseline, proving our method’s fast convergence.
We also demonstrate the advantage of the proposed U-REPA framework when measuring by parameters (rather than computation FLOPs), as shown in Fig. 5. Though U-Net brings extra parameters when FLOPs are aligned with DiTs, the advantage of SiT+U-REPA is obvious as depicted in the Parameter versus FID plot.
| ImageNet 256256, w/ cfg | ||
| Model | Epochs | FID |
| Pixel diffusion | ||
| ADM-U [9] | 400 | 3.94 |
| VDM [20] | 560 | 2.40 |
| Simple diffusion [18] | 800 | 2.77 |
| Latent Diffusion Transformer | ||
| U-ViT-H/2 [1] | 240 | 2.29 |
| DiffiT [15] | - | 1.73 |
| DiT-XL/2 [30] | 1400 | 2.27 |
| SiT-XL/2 [26] | 1400 | 2.06 |
| Masked Diffusion Transformer | ||
| MaskDiT [47] | 1600 | 2.28 |
| MDTv2-XL/2 [13] | 1080 | 1.58 |
| Representation Alignment | ||
| SiT-XL/2 + REPA [45] | 800 | 1.42 |
| SiT-XL/2 + U-REPA (Ours) | 200 | 1.48 |
| SiT-XL/2 + U-REPA (Ours) | 400 | 1.41 |
Convergence performance. We also compare our method with previous State-of-the-Arts, as shwon in Tab. 4. Our proposed SiT+U-REPA achieves a competitive FID of 1.48 with only 200 training epochs, significantly outperforming existing methods in training efficiency. Notably, while state-of-the-art masked diffusion transformers like MDTv2-XL/2 require 1,080 epochs to reach 1.58 FID, our method attains better performance (1.48) with 80% fewer iterations. Even compared to the SOTA SiT-XL/2 + REPA baseline (800 epochs for 1.42 FID), our approach uses only 1/2 of the training epochs (400) while achieving better generation quality (1.41 FID). The results demonstrate that the proposed U-REPA establishs a new efficiency frontier for diffusion models.
@@ 4.3 Ablation Studies
Encoder depths. The ablation study on encoder layer depths for feature alignment (Tab. 5) coincides with the pattern of DiT with skip connections, as we analyzed in Sec. 4: despite progressive downsampling operations that reduce spatial resolution, the centermost layers exhibit optimal alignment efficacy. For the SiT-XL/2 model, aligning features at layer 18 (midway through the 36-layer architecture) achieves peak performance with 6.25 FID and 156.2 IS, outperforming both shallower and deeper alignment points. This phenomenon persists even as the spatial dimension (Feat. Dim.) halves from 1616 to 88 in the intermediate stage, indicating that semantic richness—not spatial resolution—dominates alignment quality. Performance degradation occurs when alignment takes place at shallower or deeper stages, even though the feature size is kept the same with DINO in these stages.
@@
Feature-map upscale choices. Among the three upscaling options mentioned in Sec. 4, we figure out that upscaling U-Net hidden states after getting passed through MLP is the best option, achieving 5.72 FID and 161.6 IS. This option is also the most optimal one in terms of computation cost analysis: unlike the other two options, the downsampled size of the feature is kept throughout the whole MLP mapping process, making computation much cheaper. Overall, "Upscale after MLP" is adopted.
@@
Manifold loss weight . The ablation study on alignment weight in Eq. 5 demonstrates a clear performance peak at achieving the lowest FID (5.72) and highest IS (161.6) among tested configurations. This optimal weighting balances the alignment objective with the primary REPA loss and diffusion loss. Insufficient regularization () limits feature space coordination (FID: 6.25→5.81), while overemphasis () causes FID to regress to 5.79. Overall, we adopt in our training.

### v2

@@ Abstract
Representation Alignment (REPA) that aligns Diffusion Transformer (DiT) hidden-states with ViT visual encoders has proven highly effective in DiT training, demonstrating superior convergence properties, but it has not been validated on the canonical diffusion U-Net architecture that shows faster convergence compared to DiTs. However, adapting REPA to U-Net architectures presents unique challenges: (1) different block functionalities necessitate revised alignment strategies; (2) spatial-dimension inconsistencies emerge from U-Net’s spatial downsampling operations; (3) space gaps between U-Net and ViT hinder the effectiveness of tokenwise alignment. To encounter these challenges, we propose U-REPA, a representation alignment paradigm that bridges U-Net hidden states and ViT features as follows: Firstly, we propose via observation that due to skip connection, the middle stage of U-Net is the best alignment option. Secondly, we propose upsampling of U-Net features after passing them through MLPs. Thirdly, we observe difficulty when performing tokenwise similarity alignment, and further introduces a manifold loss that regularizes the relative similarity between samples. Experiments indicate that the resulting U-REPA could achieve excellent generation quality and greatly accelerates the convergence speed. With CFG guidance interval, U-REPA could reach in 200 epochs or 1M iterations on ImageNet 256 256, and needs only half the total epochs to perform better than REPA under sd-vae-ft-ema. Codes: https://github.com/YuchuanTian/U-REPA
@@ 1 Introduction
Representation Alignment (REPA) [repa], a methodology that aligns features from Diffusion Transformers (DiT) [dit] to modern visual encoders, has been demonstrated to significantly accelerate DiT training. This approach holds particular significance given the growing prominence of DiTs, which have gained mainstream adoption in diffusion models and are extensively applied across image generation [pixartalpha; sd3; flux] and video generation domains [opensora; opensoraplan; pyramidflow]. However, emerging empirical evidence suggests that U-Net [unet] architectures might present a more advantageous alternative to DiTs in certain scenarios [simplediffusion; hourglass; udit; dic]: U-Net-based models exhibit substantially faster convergence while achieving generation quality comparable to their transformer-based counterparts. This dichotomy motivates our core research inquiry - can modern Vision Transformer (ViT [vit])-based visual encoders be effectively adapted to guide diffusion U-Net training through alignment mechanisms similar to REPA, thereby potentially elevating the convergence speed ceiling of diffusion models?
@@ 2 Related Work
The development of diffusion architectures. The conventional diffusion works [ddpm; ddim; songunet; dhariwal] leverages a U-Net [unet] architecture, whose basic block is a concatenation of convolution layers and self-attention. More recent architectural innovations in diffusion models have witnessed a paradigm shift from conventional U-Net frameworks toward transformer-based architectures. The emergence of Diffusion Transformers [uvit; dit] demonstrates their competitive performance despite abandoning the inductive biases inherent in U-Net designs. U-ViT [uvit] represents an intermediate architecture that preserves U-Net’s hierarchical structure but replaces convolutional blocks with transformer layers, notably omitting the traditional downsampling operations. Subsequent developments have further streamlined the architecture: DiT [dit] adopts a pure transformer backbone with isotropic scaling, while SiT [sit] integrates the transformer architecture into the RectifiedFlow framework. Some other works either improve the micro-designs [visionllama; fit], or focuses on architectural efficiency [pixartsigma; sana; lit].
In contrary to these DiT works, some works still sticks to U-Net architectures and offer valuable rethinking on this conventional architectural preference: in pixel-space image generation, works including SimpleDiffusion [simplediffusion] and HourglassDiT [hourglass] still sticks to U-Net; in with its variants like U-DiT [udit], Playground v3 [playgroundv3], and DiC [dic] extending its success to latent-space diffusion through simple Conv33 designs. While these implementations empirically validate U-Net’s accelerated convergence and stable training dynamics compared to transformer-based alternatives, current research predominantly focuses on proposing architectural modifications rather than uncovering the reasons of U-Net’s superior diffusion performance.
Techniques for better DiT performance. Building upon the success of self-supervised learning [mae], MDT [mdt] and MaskDiT [maskdit] pioneer masked image modeling in diffusion frameworks by adaptively masking a good proportion of input patches during training. Other than the masking strategy, a bunch of diffusion works refer to higher-level semantic guidance from off-the-shelf pretrained models that significantly improves generation quality. REPA [repa] establishes feature alignment between ViT-based encoder embeddings and diffusion latent spaces through contrastive learning. LightningDiT [vavae] innovates through an improved VAE distilled from MAE [mae] and DINOv2 [dinov2]. Ma et al. [diffcot] introduces CLIP [clip] and DINO [dino] to verify inference-time scaling of diffusion models. These methods demonstrate that a higher-level semantic-rich feature-map from pretrained vision encoders is helpful to diffusion-based generation.
@@ 3.1 Preliminaries: REPA for DiT
Representation Alignment (REPA) [repa] distills Diffusion Transformers with semantic features from off-the-shelf ViT-based vision encoders (e.g. DINOv2 [dinov2], CLIP [clip], MAE [mae], et cetera). Given a ViT-based vision encoder and clean image , let denote its patch embeddings, where and represent the number of patches and embedding dimension, respectively. REPA establishes feature alignment between and the projected diffusion encoder outputs , where is the latent representation from the diffusion transformer at timestep , and is a trainable multilayer perceptron (MLP).
@@
where denotes a similarity metric (e.g., cosine similarity). Typically, is adopted as the output from early layers (the original work adopts layer index 8) in DiT for better alignment. This alignment term is combined with the basic flow-based diffusion objective (i.e. SiT [sit]) through a tunable coefficient , and the final loss for diffusion model training is formulated as follows:
@@ 3.2 Evaluating the Potential of U-Net
In diffusion models, U-Net and isotropic architectures (e.g., DiT) exhibit distinct design philosophies. While DiT achieves state-of-the-art results through scalability and integration with advanced techniques, U-Net-based methods emphasize faster convergence [udit]. To dissect U-Net’s efficacy, we isolate its two core components: skip connections and downsampling.
@@
Building on this insight, we propose DiT (and SiT for the flow-based version, following the naming convention of [sit]) by adding tricks of RoPE [rope] and SwiGLU following previous work [visionllama; udit]
@@ 3.3 Aligning U-Net to ViT Encoders
Position for alignment. Regarding the concern that block functionalities differ between U-Net and DiT, the comparison of prior studies [diffpruning; deltadit] demonstrates divergent hierarchical specialization: U-Net architectures typically employ mid-network layers for high-level semantic synthesis while reserving shallow layers for low-level image refinement, whereas DiTs exhibit a totally different pattern - early layers primarily govern semantic-rich outline formation with deeper layers handling detailed image refinement.
@@
Some recent works on Diffusion U-Net [unified; freeu] reveals that higher-stage U-Net features are low-frequency subspaces that discards higher frequency compoenents, including noises. Gaps are inevitable when evaluating the cosine similarities of detail-rich, high-frequency-rich vectors and flat, low-frequency-dominated vectors. In this sense, strict token-level alignment constraints like the original REPA loss prove suboptimal under these conditions, as they assume implicit feature space homogeneity between aligned modalities.
@@
Rather than strict token-wise regularization, we resort to looser objectives that does not require rigid augular alignment. Inspired by manifold knowledge distillation [manifoldKD], we hold that aligning similarities between samples from the same feature space could be a promising solution. Hence, we define Manifold Loss as
@@ 3.4 Other Improvements
We also propose and evaluate some other improvements. Due to page limits, the proposed methods and corresponding ablations are enclosed in the Appendix.
@@ 4.1 Experiment Setup
Experiment settings. Our implementation completely adheres to the training protocol established in REPA [repa]. Following the architectural configuration of latent diffusion models [sd], we employ the identical VAE variant (sd-vae-ft-ema) and adopt the AdamW optimizer. To ensure fair comparison, we maintain identical hyperparameter settings across all experiments: a global batch size of 256, fixed learning rate of , and disabled weight decay (set to 0). () is set as (0.9, 0.999). All experiments are conducted on the ImageNet 2012 benchmark [imagenet] under a controlled environment with a fixed random seed (global seed=0). 8 NVIDIA A100 GPUs are used for main experiments.
For main experiments (Tab. 8), we apply guidance interval [guidanceinterval] and SDE sampling according to the convention of REPA [repa] for fair comparison. We select smaller cfg of 1.65, because we found it is better for our architecture, different from SiT. For all ablation experiments, we train models for 100K iterations, which is sufficient to show the trend of model performance; sampling is conducted with the default setting of the official REPA codebase, i.e. in ODE and guidance interval .
@@ 4.2 The Advantage of U-REPA
Comparing SiT with SiT at different scales. We evaluate our U-REPA alignment method on ImageNet 256 under a generation setting with (REPA framework without classifier-free guidance). As shown in Table 4.2, our approach consistently improves generation quality while significantly reducing computational costs across model scales. For the base-size SiT-B/2 variant, integrating U-REPA achieves a 39.3% improvement in FID (from 24.4 to 15.3) with comparable FLOPs (24.1G vs. 23.0G) and identical training iterations (400K), demonstrating that feature alignment enhances parameter efficiency without additional training overhead. The acceleration effect becomes more pronounced in larger models: for SiT-L/2, U-REPA reduces required iterations by 42.9% (700K→400K) while simultaneously lowering FLOPs (79.3G vs. 80.8G) and achieving a 30.9% FID improvement (8.4→5.8). Most notably, the XL-scale variant with U-REPA (cf. Fig. 4.2 for FIDs vs. Training iters) attains state-of-the-art FID (5.4) using 90% fewer iterations (400K vs. 4M) and fewer FLOPs (108.8G vs. 118.6G) compared to the baseline, proving our method’s fast convergence.
We also demonstrate the advantage of the proposed U-REPA framework when measuring by parameters (rather than computation FLOPs), as shown in Fig. 4.2. Though U-Net brings extra parameters when FLOPs are aligned with DiTs, the advantage of SiT+U-REPA is obvious as depicted in the Parameter versus FID plot.
@@
Convergence performance. We also compare our method with previous State-of-the-Arts, as shown in Tab. 8. Our proposed SiT+U-REPA achieves a competitive FID of 1.48 with only 200 training epochs, significantly outperforming existing methods in training efficiency. Notably, while state-of-the-art masked diffusion transformers like MDTv2-XL/2 require 1,080 epochs to reach 1.58 FID, our method attains better performance (1.48) with 80% fewer iterations. Even compared to the SOTA SiT-XL/2 + REPA baseline (800 epochs for 1.42 FID), our approach uses only 1/2 of the training epochs (400) while achieving better generation quality (1.41 FID). The results demonstrate that the proposed U-REPA establishs a new efficiency frontier for diffusion models.
@@ 4.3 Ablation Studies
Encoder depths. The ablation study on encoder layer depths for feature alignment (Tab. 4.2) coincides with the pattern of DiT with skip connections, as we analyzed in Sec. 4: despite progressive downsampling operations that reduce spatial resolution, the centermost layers exhibit optimal alignment efficacy. For the SiT-XL/2 model, aligning features at layer 18 (midway through the 36-layer architecture) achieves peak performance with 6.25 FID and 156.2 IS, outperforming both shallower and deeper alignment points. This phenomenon persists even as the spatial dimension (Feat. Dim.) halves from 1616 to 88 in the intermediate stage, indicating that semantic richness—not spatial resolution—dominates alignment quality. Performance degradation occurs when alignment takes place at shallower or deeper stages, even though the feature size is kept the same with DINO in these stages.
| ImageNet 256256, w/ cfg | ||
| Model | Epochs | FID |
| Pixel diffusion | ||
| ADM-U [dhariwal] | 400 | 3.94 |
| VDM [vdmpp] | 560 | 2.40 |
| Simple diffusion [simplediffusion] | 800 | 2.77 |
| Latent Diffusion Transformer | ||
| U-ViT-H/2 [uvit] | 240 | 2.29 |
| DiffiT [diffit] | - | 1.73 |
| DiT-XL/2 [dit] | 1400 | 2.27 |
| SiT-XL/2 [sit] | 1400 | 2.06 |
| Masked Diffusion Transformer | ||
| MaskDiT [maskdit] | 1600 | 2.28 |
| MDTv2-XL/2 [mdt] | 1080 | 1.58 |
| Representation Alignment | ||
| SiT-XL/2 + REPA [repa] | 800 | 1.42 |
| SiT-XL/2 + U-REPA (Ours) | 200 | 1.48 |
| SiT-XL/2 + U-REPA (Ours) | 400 | 1.41 |
@@
Feature-map upscale choices. Among the three upscaling options mentioned in Sec. 4, we figure out that upscaling U-Net hidden states after getting passed through MLP is the best option, achieving 5.72 FID and 161.6 IS. This option is also the most optimal one in terms of computation cost analysis.
Manifold loss weight . The ablation study on alignment weight in Eq. 5 demonstrates a clear performance peak at achieving the lowest FID (5.72) and highest IS (161.6) among tested configurations.
4.4 Higher Resolution Experiments
At the higher-resolution ImageNet 512512 (w/ cfg) setting, U-REPA remains clearly superior to the REPA baseline (Tab. 9). Using SiT-XL/2, U-REPA reduces FID from 2.44 to 2.21 and raises IS from 247.3 to 274.7. These results indicate that U-REPA’s benefits persist at 512 resolution, yielding better distributional fidelity and sample quality/diversity, and demonstrating strong scalability.
| ImageNet 512512, w/ cfg | ||
| Alignment Choices | FID | IS |
| SiT-XL/2 + REPA | 2.44 | 247.3 |
| SiT-XL/2 + U-REPA (Ours) | 2.21 | 274.7 |
4.5 The Energy Cost Advantage of U-REPA
We also assess the energy-cost advantage of U-REPA over REPA. We train on eight NVIDIA A100 GPUs and record each GPU’s power draw. Combining the measured power with the training duration, we estimate the total energy consumed. The statistics for average power and estimated total energy used by all 8 GPUs are summarized in Tab. 10. Results indicate that our U-REPA method is "greener", costing far less energy.
Reducing training energy directly curbs operational emissions. Methods that achieve comparable accuracy with lower energy, such as U-REPA vs. REPA in our study, advance both sustainability and the economic viability of large-scale AI.
| ImageNet 256256 | ||||
| Model | Avg. Pow. (W) | Training Hours | Est. Energy (J) | FID |
| SiT-XL/2+REPA (4M iter) | 373.2 | 302.7 | 3.2510e9 | 1.42 |
| SiT-XL/2+U-REPA (2M iter) | 295.3 | 230.7 | 1.9610e9 | 1.41 |